OpenStudy (anonymous):

A sandbag was thrown downward from a building. The function f(t) = -16t2 - 32t + 384 shows the height f(t), in feet, of the sandbag after t seconds: Part A: Factor the function f(t) and use the factors to interpret the meaning of the x-intercept of the function. Part B: Complete the square of the expression for f(x) to determine the vertex of the graph of f(x). Would this be a maximum or minimum on the graph? Part C: Use your answer in part B to determine the axis of symmetry for f(x)?

OpenStudy (anonymous):

for part A i got -16(t+6)(x-4) when factored and said that the x-intercepts of the functions is (4,0) which represents the time in seconds that the sandbag was thrown to where it lands.

OpenStudy (anonymous):

For part b I completed the square and got (t+1)^2 = 25 but how do I find the vertex?

OpenStudy (anonymous):

@Hero @Hero you said you would help me!

hero (hero):

I said, post your next question separately. Anyone can help with these.

OpenStudy (anonymous):

Yeah I did post it separately and you said you would help!

OpenStudy (anonymous):

I need help on finding the vertex and I am confused on how to

TheSmartOne (thesmartone):

For Part A, you are correct (almost)

OpenStudy (anonymous):

What did I leave out? @TheSmartOne could you please explain

TheSmartOne (thesmartone):

It's a simply mistake, but: You were given \(\sf\Large f(t) = -16t^2 - 32t + 384\) however you factored it to \(\sf\large -16(t+6)(\color{red}{x}-4)\)

OpenStudy (anonymous):

so the t in t-4 is supposed to be x-4? @TheSmartOne

TheSmartOne (thesmartone):

The variable you were given @kaite_mcgowan was \(\bf t\) so I don't understand where you brought in the \(\bf x\) from. So you just need to change the x to a t to make it correct :)

OpenStudy (anonymous):

oh ok thanks so much! could you please help me on b and finding the vertex with the completed square of (t+1)^2 = 25 @TheSmartOne

TheSmartOne (thesmartone):

\(\sf\Large y=a(x-h)^2+k\) The vertex is \(\sf\Large (h,k)\)

OpenStudy (anonymous):

so how would I incorporate (t+1)^2 = 25 in to that?

OpenStudy (anonymous):

y = a(t-1)^2 + 25 ???? @TheSmartOne

TheSmartOne (thesmartone):

one sec

TheSmartOne (thesmartone):

You didn't properly complete the square.

OpenStudy (anonymous):

?? what do you mean?

TheSmartOne (thesmartone):

\(\sf\Large -16t^2 - 32t + 384 \color{red}{\neq} (t-1)^2+25\)

OpenStudy (anonymous):

so the completed square form is (t - 1)^2 + 25

TheSmartOne (thesmartone):

it isn't. That's wrong.

OpenStudy (anonymous):

oh ok i got it. How would I find it then?

OpenStudy (anonymous):

could you show me step by step

TheSmartOne (thesmartone):

\(\sf\Large -16t^2 - 32t =-384\) What is \(\sf\Large\left( \frac{b}{2}\right)^2=\left( \frac{-32}{2}\right)^2=?\)

OpenStudy (anonymous):

-16

TheSmartOne (thesmartone):

correct; (-16)^2

OpenStudy (anonymous):

whats next ??? @TheSmartOne

TheSmartOne (thesmartone):

you have to add it to both sides: \(\sf\Large -16t^2 - 32t+(-16^2) =-384+(-16)^2\)

TheSmartOne (thesmartone):

can you factor the left hand side? Hint: \(\sf\Large a^2+2ab+b^2=(a+b)^2\)

OpenStudy (anonymous):

I am so sorry i lost wifi! @TheSmartOne

OpenStudy (anonymous):

ok so you get -16t^2 - 32t + (-256) = -384 + (-256) @TheSmartOne

OpenStudy (anonymous):

-16t^2 - 32t + (-256) = -640 @TheSmartOne

TheSmartOne (thesmartone):

hmmm

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

One sec, we'll need to backtrack over here

OpenStudy (anonymous):

ok no problem

OpenStudy (anonymous):

do you know what went wrong?

TheSmartOne (thesmartone):

we have to first factor the 16 out of the equation, so: \(\sf\Large -16t^2 - 32t + 384 =-16(t^2+2t-24) \)

TheSmartOne (thesmartone):

@Mehek14 @paki I'm not sure how to complete the square :/

OpenStudy (anonymous):

so the factored form is -16(t+6)(t-4)

TheSmartOne (thesmartone):

there are so many ways to solve math, I don't know why they limit us with this completing the square lol

Mehek (mehek14):

*disappears*

OpenStudy (anonymous):

wait thats the answer the link above? I know that the highest point or the maximum is 400 and the symmetry of axis is -1???

OpenStudy (anonymous):

i just dont understand completing the square?

TheSmartOne (thesmartone):

the link was an example of how to do it

OpenStudy (anonymous):

oh ok so what would I do ??

TheSmartOne (thesmartone):

yes, the symmetry of axis is x=-1

OpenStudy (anonymous):

ok I go that part and I know that the maximum is 400 i just need to show my work of how I got that

TheSmartOne (thesmartone):

@zepdrix could you help us to complete the square for \(\sf\Large f(t)=-16t^2-32t+384\)

OpenStudy (anonymous):

@zepdrix please I really need help!

Mehek (mehek14):

I can tell you how to do the first step

Mehek (mehek14):

divide all the numbers by -16

TheSmartOne (thesmartone):

the calculator gives you the final answer, but no steps: http://prntscr.com/7metrk

TheSmartOne (thesmartone):

@mathmate could you help us complete the square?

OpenStudy (anonymous):

t^2 + 2t +24

OpenStudy (anonymous):

OMG EVERYONE I THINK I MAY HAVE FIGURED IT OUT! GIVE ME ONE SEC!

TheSmartOne (thesmartone):

we could have just as easily found the vertex by -b/2a ¯\_(ツ)_/¯

OpenStudy (anonymous):

a(x+d)2+e d=−32 / 2⋅(−16) d=1 e=384− (−32)^2 / 4⋅(−16) e=400 −16(t+1)2+400

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (jtvatsim):

that's it, I've been working it out over here as well. nicely done. :)

OpenStudy (anonymous):

thanks so much!!!! @jtvatsim

TheSmartOne (thesmartone):

well we got the completing square out of the way

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @TheSmartOne \(\sf\Large y=a(x-h)^2+k\) The vertex is \(\sf\Large (h,k)\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

haha yep! so to find the vertex I wold just do y = -16(x-1)^2 +400??

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

or in other words the vertex is (-1,400)

OpenStudy (anonymous):

y = -16(x+1)^2 +400

TheSmartOne (thesmartone):

correct

TheSmartOne (thesmartone):

and that would be a maximum or a minimum?

OpenStudy (anonymous):

maximum!

TheSmartOne (thesmartone):

correct :)

OpenStudy (anonymous):

Thanks for all of your help @TheSmartOne and sticking with me through the whole equation unlike other people who promised before **cough cough @Hero cough cough** thanks again @TheSmartOne

hero (hero):

I had to go to lunch. I would have helped otherwise.

OpenStudy (anonymous):

yeah yeah whatever

TheSmartOne (thesmartone):

I probably wouldn't have helped if it wasn't for Hero: http://prntscr.com/7mezsj

OpenStudy (anonymous):

at least hero considered it...

OpenStudy (mathmate):

@kaite_mcgowan Is the question resolved?

OpenStudy (anonymous):

yes it is thanks for asking! @mathmate

OpenStudy (mathmate):

Sorry, I came a little late. Here's what I would have done anyway: Given f(t)=\(-16t^2-32t+384\) (a) factorization f(t)=\(-16(t^2+2t-24)\) f(t)=-16(t+6)(t-4) (b) Complete square f(t)=-16(t+1 +5)(t+1 -5) f(t)=-16[(t+1)^2-25] maximum is at x=-1 (solve for t in t+1=0), f(-1)=400. Maximum because leading coefficient (of t^2) is negative. (c) Axis of symmetry Axis of symmetry of a quadratic is location of maximum/minimum, i.e. x=-1